OpenStudy (rootbeer003):

What is the cosine of a 30° angle?

OpenStudy (wolf1728):

Do you have a book with a list of cosine values? Do you have a calculator with the cosine function?

OpenStudy (rootbeer003):

umm

OpenStudy (rootbeer003):

I have the TI-83 plus?

OpenStudy (wolf1728):

That might have the cosine function. (I'm not familiar with the TI-83)

OpenStudy (rootbeer003):

Ya it does

OpenStudy (wolf1728):

Try using it

OpenStudy (rootbeer003):

My answer has to be a sq root fraction

OpenStudy (wolf1728):

Hmmm, maybe they want you to actually calculate it. I can help you.

OpenStudy (rootbeer003):

The calculator on here has it

OpenStudy (rootbeer003):

o nvm

OpenStudy (rootbeer003):

I think its sq root 3

OpenStudy (wolf1728):

Here is a graphic

OpenStudy (rootbeer003):

1/2

OpenStudy (wolf1728):

That is a 30-60-90 triangle. In such a triangle the hypotenuse is exactly twice as big as the small side.

OpenStudy (rootbeer003):

oo

OpenStudy (wolf1728):

So hypotenuse = 2 Opposite = 1 Then we need the Pythagoean Theorem to find the adjacent side 2^2 = 1^2 + adj^2 4 - 1 = adjacent side squared 3 = adjacent side squared adjacent side = square root of 3 (we're not done yet)

OpenStudy (wolf1728):

Looking at the graphic, we see that the cosine of 30 degrees = adjacent side / hypotenuse cosine of 30 degrees = (square root of 3) / 2 and there's your answer It's that simple -- LOL

OpenStudy (rootbeer003):

Ty, can u help me with 3 more?

OpenStudy (wolf1728):

Well let's try 1 more :-)

OpenStudy (rootbeer003):

What is the cosine of a 30° angle?

OpenStudy (wolf1728):

You've already asked that

OpenStudy (rootbeer003):

o

OpenStudy (rootbeer003):

i meant tan

OpenStudy (wolf1728):

Okay tangent of 30 degrees Look at the graphic

OpenStudy (wolf1728):

We know from the previous problem that the opposite side = 1 and the adjacent side = square root of 3 Therefore the tangent of 30 degrees = 1/(square of 3)

OpenStudy (rootbeer003):

uh thats not a choice

OpenStudy (wolf1728):

What are the choices?

OpenStudy (wolf1728):

That equals the sq root of 3 / 3 (Don't know how they got that but it's correct (SO is 1/sq root(3)

OpenStudy (rootbeer003):

mk

OpenStudy (wolf1728):

Multiply 1/(sq root of 3) by the sq root of 3 and you get sq root (3) / 3 :-)

OpenStudy (rootbeer003):

ooooo

OpenStudy (rootbeer003):

So can u help me with 2 more? or check my answers?

OpenStudy (wolf1728):

Sure - go ahead

OpenStudy (rootbeer003):

What is the tangent of a 60° angle? I put sq root 3

OpenStudy (wolf1728):

Well I made another graphic

OpenStudy (rootbeer003):

is it sq root 3/2?

OpenStudy (wolf1728):

hyp = 2 adj = 1 from Pythagorean Theorem hyp^2 = adj^2 + opp^2 opp^2 = hyp^2 - adj^2 opp^2 = 4 -1 opp^2 = 3 opposite side = sq root of 3 tan 60 = opp/adj tan 60 = sq root of 3 / 1 = sq root of 3 Very good Root beer

OpenStudy (wolf1728):

No wait it's just square root of 3

OpenStudy (rootbeer003):

o ok haha

OpenStudy (wolf1728):

You were right the first time

OpenStudy (rootbeer003):

What is the sine of a 30° angle I puy sq root 3/3

OpenStudy (rootbeer003):

put*

OpenStudy (wolf1728):

Here's the new graphic Remember the hypotenuse is 2 and the opposite side = 1 So, the sine of 30 degrees equals??

OpenStudy (rootbeer003):

1/2?

OpenStudy (wolf1728):

That's right By the way, here's an online calculator if you want to check the answers (and they are correct) http://www.1728.org/trigcalc.htm

OpenStudy (wolf1728):

Incidentally, that's my website :-)

OpenStudy (rootbeer003):

lol ok ty

OpenStudy (wolf1728):

u r welcome May I have a medal?

OpenStudy (wolf1728):

Thanks :-)

OpenStudy (rootbeer003):

np